# Scalable and Robust Multiband Modeling of AGN Light Curves in Rubin-LSST

### v1

@@ Abstract
software: pandas (McKinney, 2010), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020), matplotlib (Hunter, 2007), astropy (The Astropy Collaboration et al., 2013; The Astropy Collaboration et al., 2018; Astropy Collaboration et al., 2022), emcee (Foreman-Mackey et al., 2013), eztao (Yu & Richards, 2022), tinygp (Foreman-Mackey et al., 2024), numpyro (Bingham et al., 2019; Phan et al., 2019) eztaox (Yu & Burke, 2025)
@@ 1 Motivation
The UV/optical continuum luminosity of active galactic nuclei (AGNs) exhibit correlated stochastic variability on timescales ranging from hours to years (e.g., Vanden Berk et al., 2004; Sesar et al., 2007; Schmidt et al., 2012). This variability is believed to originate in the accretion disk, potentially driven in part by thermal reprocessing of X-ray photons emitted from the central corona (Krolik et al., 1991) and/or accretion rate fluctuations propagating through the disk (Lyubarskii, 1997). For a recent review of the observational properties of AGN variability in the UV/optical and X-ray bands, we refer the reader to Paolillo & Papadakis (2025).
@@
A damped random walk (DRW) process is commonly adopted to model AGN UV/optical light curves (Kelly et al., 2009). The power spectral density (PSD) of a DRW exhibits a broken power-law (PL; $1/f^{\alpha}$) form, where the PL exponent $\alpha$ transitions from zero at low frequencies to two at high frequencies. The point at which this change in PL exponent occurs is known as the break frequency, which corresponds to a characteristic timescale denoted by $\tau_{\rm DRW}$. The asymptotic variability amplitude of the process is denoted by $\sigma_{\mathrm{DRW}}$. Although the DRW process is a phenomenological model and the precise PSD shape of intrinsic AGN UV/optical variability remains under debate (e.g., Mushotzky et al., 2011; Kasliwal et al., 2015; Smith et al., 2018; Arévalo et al., 2024; Beard et al., 2025; Petrecca et al., 2024; Yu et al., 2025), it continues to be a valuable and widely used tool for modeling AGN light curves because of its adequacy and computational efficiency.
@@
Modeling AGN UV/optical light curves as a DRW has revealed compelling correlations between variability and AGN fundamental properties. Numerous previous investigations have shown that the best-fit $\sigma_{\mathrm{DRW}}$ correlates negatively with the Eddington ratio ($L/L_{\rm Edd}$) and bolometric luminosity ($L_{\rm bol}$) of AGNs (e.g., MacLeod et al., 2010; Sánchez-Sáez et al., 2018; Suberlak et al., 2021). Additionally, the characteristic timescale of the DRW model ($\tau_{\rm DRW}$) correlates positively with $M_{\rm{BH}}$ (e.g., Burke et al., 2021; Wang et al., 2023). These empirical relations not only provide valuable insights into the physical mechanisms driving AGN variability, but also form the basis for new techniques to estimate AGN fundamental properties using photometric data alone. For example, Burke et al. (2022) leveraged the $\tau_{\rm DRW}$$M_{\rm{BH}}$ correlation to search for low-mass AGNs powered by accreting massive black holes (MBH; $10^{5}\,$$M_{\odot}$$M_{\rm{BH}}$$10^{7}\,$$M_{\odot}$) at the centers of dwarf galaxies  (i.e., $M_{\star}\lesssim 10^{10}\,M_{\odot}$; Greene et al., 2020).
@@
AGN UV/optical variability at longer wavelengths is also often observed to lag behind the variability at shorter wavelengths (Collier et al., 1999; Sergeev et al., 2005). Under the thermal reprocessing scenario, longer-wavelength photons originate from disk annuli located farther from the central supermassive black hole (SMBH) than those emitting shorter-wavelength photons (Shakura & Sunyaev, 1973). Consequently, the reprocessed emission at longer wavelengths exhibits longer light-crossing time delays relative to the central irradiating source (Cackett et al., 2007). This specific geometry—in which a compact source near the SMBH (e.g., an X-ray corona) irradiates the accretion disk—is commonly referred to as the lamp-post model. In this framework, the size of the accretion disk can be inferred by measuring the time lags between UV/optical light curves at different wavelengths—a technique known as continuum reverberation mapping (CRM), which probes the geometry of unresolved sources through time-delay measurements (see Cackett et al., 2021, and references therein).
@@
Numerous CRM experiments have reported that AGN accretion disk sizes are systematically larger than those predicted by the standard thin-disk model (e.g., Fausnaugh et al., 2016; Jiang et al., 2017; Jha et al., 2022). Larger-than-expected accretion disk sizes have also been independently inferred from microlensing observations of gravitationally lensed quasars, where the disk half-light radius (i.e., the characteristic size) is inferred from the differential microlensing variability exhibited by the multiple lensed images (e.g., Pooley et al., 2007; Dai et al., 2010; Morgan et al., 2010; Jiménez-Vicente et al., 2014; Cornachione et al., 2020). It has been suggested that the measured lags may be contaminated by reprocessed UV/optical photons originating in the broad emission-line region (BLR), which lies farther out in the accretion flow (e.g., Korista & Goad, 2001; McHardy et al., 2018; Cackett et al., 2018; Lawther et al., 2018). This contamination could cause the observed interband continuum lags to appear longer than those expected from pure disk reprocessing.
@@
In contrast, some studies have argued that accretion disk sizes inferred from CRM are consistent with the standard thin-disk model when factors such as temperature fluctuations within the disk or potential selection biases are properly accounted for (e.g., Dexter & Agol, 2011; Mudd et al., 2018; Homayouni et al., 2019; Yu et al., 2020b). Moreover, Kammoun et al. (2021) showed that the standard thin-disk model can reproduce the observed CRM lags when relativistic (special and general) and disk ionization effects are fully included. The accretion disk size discrepancy poses one of the significant challenges to the standard thin-disk model and to our understanding of AGN accretion physics more broadly. A large sample of AGNs with robustly measured CRM lags is required to investigate this issue further. To date, however, CRM lags have been successfully measured for only a few hundred AGNs (e.g., Collier et al., 1998; Peterson et al., 1998; Sergeev et al., 2005; Jiang et al., 2017; Homayouni et al., 2019; Mudd et al., 2018; Yu et al., 2020b; Guo et al., 2022).
@@
The high-quality light curves provided by LSST will enable robust characterization of AGN UV/optical variability and interband continuum lags. Interband lags can be measured for about one million LSST AGNs (Yu et al., 2020a, Li et al. 2025 submitted), allowing detailed investigations of the accretion disk size problem as a function of AGN properties such as $L_{\rm Edd}$ and $M_{\rm BH}$, and offering new insights into the physical origin of continuum lags. Meanwhile, stochastic variability parameters—such as characteristic timescales and amplitudes—extracted from LSST light curves will enable reliable AGN classification (e.g., Kozłowski et al., 2010; Butler & Bloom, 2011; MacLeod et al., 2011; Peters et al., 2015; De Cicco et al., 2021; Sánchez-Sáez et al., 2021; Savić et al., 2023; Sánchez-Sáez et al., 2023). For example, variability-based methods are more effective than SED-based methods at identifying low-luminosity and low-mass AGNs (e.g., in dwarf galaxies), where host-galaxy starlight contamination is significant (e.g., Baldassare et al., 2017; Ward et al., 2022; Bernal et al., 2025). A complete census of low-luminosity is essential for understanding AGN evolution across cosmic time (Richards et al., 2006), while a comprehensive inventory of low-mass AGNs is critical for constraining the formation pathways of SMBHs (see Greene et al., 2020, and references therein).
However, existing AGN light curve analysis tools are not well-prepared to fully exploit the scale and richness of LSST observations. Current tools/methods used to characterize AGN UV/optical variability typically operate on single-band light curves and lack the ability to incorporate multiband information (e.g., Vaughan et al., 2003; Kelly et al., 2009, 2014; Kozłowski, 2016; Yu & Richards, 2022). This limitation prevents the effective utilization of LSST’s multiband coverage for improved characterization of AGN variability. Tools/methods commonly used to measure interband lags generally fall into two categories. The first includes techniques that estimate lags by computing the cross-correlation function (CCF) between light curves; the CCF reaches its maximum when one light curve is shifted relative to the other by an amount corresponding to the intrinsic lag (e.g., Gaskell & Peterson, 1987; Edelson & Krolik, 1988; Alexander, 1997). [^1] When applied to a single-band light curve and its shifted copies, the CCF method yields its autocorrelation function (ACF), which can be used to characterize AGN variability (e.g., Raiteri et al., 2021a, b). These methods are typically computationally efficient, but are limited to lag estimation and do not provide a full characterization of the underlying variability process. The second consists of methods that model each observed light curve as the convolution of an unknown driving continuum with a time-lag distribution, represented by a transfer function (e.g., Zu et al., 2011; Li et al., 2016; Starkey et al., 2016; Donnan et al., 2021). This is expressed as,
@@
where $y_{\rm drive}(t-\tau)$ denotes the flux of the underlying driving continuum at $t-\tau$, and $\Psi(\tau)$ represents the corresponding transfer function. JAVELIN is a representative tool in this category (Zu et al., 2013). It models the driving continuum as a DRW process and adopts a top-hat profile for the transfer function. While such methods provide a more comprehensive treatment by jointly modeling variability and interband lags, they are often computationally demanding.
Many recent investigations have turned to machine learning (ML) techniques to develop more scalable methods for analyzing AGN light curves (e.g., Fagin et al., 2024; Li et al., 2024). In these approaches, deep neural networks are trained on simulated light curves and are tasked with recovering the underlying variability characteristics and/or the physical parameters used to generate the simulations. However, the reliability of such methods hinges on the assumption that the models used to generate the training data accurately represent both the variability physics and the observational conditions (e.g., cadence, photometric uncertainty)—an assumption rarely satisfied in practice. For example, Li et al. (2024) show that the AGN physical parameters (e.g., $M_{\rm{BH}}$) predicted by their simulation-based inference model (Cranmer et al., 2020) can be biased by up to 0.5 dex when the test light curves are five times noisier than those used for training.
In this paper, we present EzTaoX—a scalable and flexible AGN light curve modeling tool that leverages LSST’s multiband observations to simultaneously infer the properties of the underlying stochastic variability and the lags between photometric bands (Yu & Burke, 2025). The name of EzTaoX combines three elements: “Ez", a homophone of easy; “Tao" referencing the Chinese philosophical and religious tradition of Taoism, commonly translated as “the Way"; and “X" denoting that the code is implemented in JAX—a high-performance numerical computing library developed by Google. EzTaoX achieves a speed increase of $\sim 10^{2}-10^{4}\times$ on CPUs compared to existing tools with similar capabilities (e.g., JAVELIN), making it feasible to perform both stochastic variability characterization and interband continuum lag measurements for all AGNs discovered by LSST. Furthermore, EzTaoX’s multiband capability enables robust constraints on stochastic variability timescales (e.g., $\tau_{\rm DRW}$) down to one hour with the LSST WFD cadence, facilitating the identification of low-mass AGNs—such as those residing in dwarf galaxies—which are expected to exhibit characteristic $\tau_{\rm DRW}$ ranging from hours to days (Burke et al., 2021; Wang et al., 2023).
@@ 2 Light Curve Modeling Method & Implementation
EzTaoX models AGN UV/optical stochastic variability as a Gaussian processes (GP), a method commonly adopted in AGN variability analysis (e.g., Press et al., 1992; Kelly et al., 2009; Kozłowski et al., 2010; Wilkins, 2019; Griffiths et al., 2021; Yu et al., 2022; Stone et al., 2022; Zhang et al., 2023; McLaughlin et al., 2024). Since the flux light curves of accreting compact objects typically follow a lognormal distribution (e.g., Uttley & McHardy, 2001; Gaskell, 2004; Giebels & Degrange, 2009; MacLeod et al., 2012; De Cicco et al., 2022), we emphasize that GP modeling is best applied to magnitude light curves, which follow a Gaussian distribution. [^2] Gúrpide & Middleton (2025) demonstrated that even when the light curves follow a lognormal distribution, the simulated GP parameters can be recovered without significant bias.
Given a single band AGN light curve, a GP treats the intrinsic brightness of the AGN at each point in time as a Gaussian random variable and captures the temporal correlations between observations through a kernel function $k(t_{i},t_{j})$, which defines the covariance between the Gaussian variables at any two timestamps $t_{i}$ and $t_{j}$ (Rasmussen & Williams, 2006; Aigrain & Foreman-Mackey, 2023).
@@
where $k_{\rm latent}$ denotes the kernel of the underlying latent GP. $S_{1}$ and $S_{2}$ are the variability-amplitude scaling factors for the corresponding photometric bands, and $\tau_{\rm lag}$ represents the relative time lag between them. The variability amplitude in each band is obtained by multiplying the latent GP amplitude by the corresponding scaling factor. This class of kernel functions, as expressed in Equation 5, has seen broad application in astronomical time series analysis (e.g., Gordon et al., 2020; Villar et al., 2021). To evaluate the likelihood of the multiband model for a given light curve, we shift the light curve in different bands by their trial $\tau_{\rm lag}$ relative to the reference band, and compute Equation 2 using a kernel that assumes no intrinsic lag between bands (i.e., setting $\tau_{\rm lag}=0$ in Equation 5). In this way, the GP parameters and the interband lags are fitted jointly.
@@
EzTaoX is built on top of tinygp [^3] https://github.com/dfm/tinygp—a GP modeling framework developed using JAX. [^4] https://github.com/jax-ml/jax JAX is a high-performance numerical computing library that combines automatic differentiation, just-in-time (JIT) compilation, and hardware acceleration to empower scalable machine learning and scientific computing. tinygp leverages these features of JAX and re-implements the novel celerite algorithm introduced by Foreman-Mackey et al. (2017) in JAX. When a kernel function can be expressed as a mixture of exponentials, the celerite algorithm can evaluate the GP likelihood function with $\mathcal{O}(N)$ computational scaling, a significant efficiency enhancement over the standard $\mathcal{O}(N^{3})$ scaling associated with direct likelihood evaluation (Foreman-Mackey et al., 2017). EzTaoX takes advantage of the JAX-based implementation of the celerite algorithm to provide a scalable solution for multiband light curve modeling of AGNs.
@@
PSD shapes exhibited by the light curves of accreting black holes can be effectively modeled with celerite GP kernels. For example, the continuous-time autoregressive moving-average (CARMA) process—commonly used as a more flexible alternative to the DRW process—has an equivalent celerite implementation (Roux, 2002; Kelly et al., 2014; Kasliwal et al., 2017; Moreno et al., 2019; Yu et al., 2022). A CARMA($p$, $q$) process combines an autoregressive process of order $p$ with a moving-average process of order $q$, and its PSD can be expressed as a weighted sum of modified Lorentzians. The DRW process is a special case of a CARMA process with $p=1$ and $q=0$. For a detailed description of CARMA modeling, we refer the reader to Kelly et al. (2014). More recently, Lefkir et al. (2025) introduced an algorithm for approximating arbitrary broken PL PSDs using a sum of multiple celerite kernels, thereby enhancing the interpretability and flexibility of GP-based modeling for AGN light curves. Moreover, celerite-based kernels can be freely combined through addition and multiplication to model more complex covariance structures, while retaining the $\mathcal{O}(N)$ computational scaling. All such kernels can be used for simultaneous multiband variability characterization and interband lag measurements.

### v2

@@ Abstract
software: pandas (pandas2010), numpy (numpy2020), scipy (scipy2020), matplotlib (matplotlib2007), astropy (astropy2013; theastropycollaboration2018; astropycollaboration2022), emcee (foreman-mackey2013), eztao (yu2022), tinygp (foreman-mackey2024), numpyro (bingham2019pyro; phan2019composable) eztaox (eztaox)
@@ 1 Motivation
The UV/optical continuum luminosity of active galactic nuclei (AGNs) exhibit correlated stochastic variability on timescales ranging from hours to years (e.g., vandenberk2004; sesar2007; schmidt2012). This variability is believed to originate in the accretion disk, potentially driven in part by thermal reprocessing of X-ray photons emitted from the central corona (Krolik1991) and/or accretion rate fluctuations propagating through the disk (lyubarskii1997). For a recent review of the observational properties of AGN variability in the UV/optical and X-ray bands, we refer the reader to Paolillo2025.
@@
A damped random walk (DRW) process is commonly adopted to model AGN UV/optical light curves (kelly2009). The power spectral density (PSD) of a DRW exhibits a broken power-law (PL; $1/f^{\alpha}$) form, where the PL exponent $\alpha$ transitions from zero at low frequencies to two at high frequencies. The point at which this change in PL exponent occurs is known as the break frequency, which corresponds to a characteristic timescale denoted by $\tau_{\rm DRW}$. The asymptotic variability amplitude of the process is denoted by $\sigma_{\mathrm{DRW}}$. Although the DRW process is a phenomenological model and the precise PSD shape of intrinsic AGN UV/optical variability remains under debate (e.g., mushotzky2011; kasliwal2015; smith2018; Arevalo2024; Beard2025; Petrecca2024; Yu2025), it continues to be a valuable and widely used tool for modeling AGN light curves because of its adequacy and computational efficiency.
@@
Modeling AGN UV/optical light curves as a DRW has revealed compelling correlations between variability and AGN fundamental properties. Numerous previous investigations have shown that the best-fit $\sigma_{\mathrm{DRW}}$ correlates negatively with the Eddington ratio ($L/L_{\rm Edd}$) and bolometric luminosity ($L_{\rm bol}$) of AGNs (e.g., macleod2010; sanchez-saez2018; suberlak2021). Additionally, the characteristic timescale of the DRW model ($\tau_{\rm DRW}$) correlates positively with $M_{\rm{BH}}$ (e.g., burke2021; Wang2023a). These empirical relations not only provide valuable insights into the physical mechanisms driving AGN variability, but also form the basis for new techniques to estimate AGN fundamental properties using photometric data alone. For example, burke2022 leveraged the $\tau_{\rm DRW}$$M_{\rm{BH}}$ correlation to search for low-mass AGNs powered by accreting massive black holes (MBH; $10^{5}\,$$M_{\odot}$$M_{\rm{BH}}$$10^{7}\,$$M_{\odot}$) at the centers of dwarf galaxies  (i.e., $M_{\star}\lesssim 10^{10}\,M_{\odot}$; greene2020).
@@
AGN UV/optical variability at longer wavelengths is also often observed to lag behind the variability at shorter wavelengths (Collier1999; sergeev2005). Under the thermal reprocessing scenario, longer-wavelength photons originate from disk annuli located farther from the central supermassive black hole (SMBH) than those emitting shorter-wavelength photons (shakura1973). Consequently, the reprocessed emission at longer wavelengths exhibits longer light-crossing time delays relative to the central irradiating source (Cackett2007). This specific geometry—in which a compact source near the SMBH (e.g., an X-ray corona) irradiates the accretion disk—is commonly referred to as the lamp-post model. In this framework, the size of the accretion disk can be inferred by measuring the time lags between UV/optical light curves at different wavelengths—a technique known as continuum reverberation mapping (CRM), which probes the geometry of unresolved sources through time-delay measurements (see cackett2021, and references therein).
@@
Numerous CRM experiments have reported that AGN accretion disk sizes are systematically larger than those predicted by the standard thin-disk model (e.g., fausnaugh2016; jiang2017; Jha2022). Larger-than-expected accretion disk sizes have also been independently inferred from microlensing observations of gravitationally lensed quasars, where the disk half-light radius (i.e., the characteristic size) is inferred from the differential microlensing variability exhibited by the multiple lensed images (e.g., Pooley2007; Dai2010; morgan2010; Jimenez-Vicente2014; Cornachione2020). It has been suggested that the measured lags may be contaminated by reprocessed UV/optical photons originating in the broad emission-line region (BLR), which lies farther out in the accretion flow (e.g., korista2001; mchardy2018; cackett2018; lawther2018). This contamination could cause the observed interband continuum lags to appear longer than those expected from pure disk reprocessing.
@@
In contrast, some studies have argued that accretion disk sizes inferred from CRM are consistent with the standard thin-disk model when factors such as temperature fluctuations within the disk or potential selection biases are properly accounted for (e.g., dexter2011a; mudd2018; homayouni2019; yu2020). Moreover, Kammoun2021 showed that the standard thin-disk model can reproduce the observed CRM lags when relativistic (special and general) and disk ionization effects are fully included. The accretion disk size discrepancy poses one of the significant challenges to the standard thin-disk model and to our understanding of AGN accretion physics more broadly. A large sample of AGNs with robustly measured CRM lags is required to investigate this issue further. To date, however, CRM lags have been successfully measured for only a few hundred AGNs (e.g., Collier1998; Peterson1998; sergeev2005; jiang2017; homayouni2019; mudd2018; yu2020; guo2022).
@@
The high-quality light curves provided by LSST will enable robust characterization of AGN UV/optical variability and interband continuum lags. Interband lags can be measured for about one million LSST AGNs (yu2020a, Li et al. 2025 submitted), allowing detailed investigations of the accretion disk size problem as a function of AGN properties such as $L_{\rm Edd}$ and $M_{\rm BH}$, and offering new insights into the physical origin of continuum lags. Meanwhile, stochastic variability parameters—such as characteristic timescales and amplitudes—extracted from LSST light curves will enable reliable AGN classification (e.g., kozlowski2010; butler2011; macleod2011; peters2015; DeCicco2021; Sanchez-Saez2021; savic2023; Sanchez-Saez2023). For example, variability-based methods are more effective than SED-based methods at identifying low-luminosity and low-mass AGNs (e.g., in dwarf galaxies), where host-galaxy starlight contamination is significant (e.g., baldassare2017; ward2022; bernal2025). A complete census of low-luminosity is essential for understanding AGN evolution across cosmic time (richards2006a), while a comprehensive inventory of low-mass AGNs is critical for constraining the formation pathways of SMBHs (see greene2020, and references therein).
However, existing AGN light curve analysis tools are not well-prepared to fully exploit the scale and richness of LSST observations. Current tools/methods used to characterize AGN UV/optical variability typically operate on single-band light curves and lack the ability to incorporate multiband information (e.g., Vaughan2003; kelly2009; kelly2014; kozlowski2016; yu2022). This limitation prevents the effective utilization of LSST’s multiband coverage for improved characterization of AGN variability. Tools/methods commonly used to measure interband lags generally fall into two categories. The first includes techniques that estimate lags by computing the cross-correlation function (CCF) between light curves; the CCF reaches its maximum when one light curve is shifted relative to the other by an amount corresponding to the intrinsic lag (e.g., gaskell1987; edelson1988; Alexander1997). [^1] When applied to a single-band light curve and its shifted copies, the CCF method yields its autocorrelation function (ACF), which can be used to characterize AGN variability (e.g., Raiteri2021a; Raiteri2021b). These methods are typically computationally efficient, but are limited to lag estimation and do not provide a full characterization of the underlying variability process. The second consists of methods that model each observed light curve as the convolution of an unknown driving continuum with a time-lag distribution, represented by a transfer function (e.g., zu2011; li2016; starkey2016; donnan2021). This is expressed as,
@@
where $y_{\rm drive}(t-\tau)$ denotes the flux of the underlying driving continuum at $t-\tau$, and $\Psi(\tau)$ represents the corresponding transfer function. JAVELIN is a representative tool in this category (zu2013). It models the driving continuum as a DRW process and adopts a top-hat profile for the transfer function. While such methods provide a more comprehensive treatment by jointly modeling variability and interband lags, they are often computationally demanding.
Many recent investigations have turned to machine learning (ML) techniques to develop more scalable methods for analyzing AGN light curves (e.g., fagin2024; li2024). In these approaches, deep neural networks are trained on simulated light curves and are tasked with recovering the underlying variability characteristics and/or the physical parameters used to generate the simulations. However, the reliability of such methods hinges on the assumption that the models used to generate the training data accurately represent both the variability physics and the observational conditions (e.g., cadence, photometric uncertainty)—an assumption rarely satisfied in practice. For example, li2024 show that the AGN physical parameters (e.g., $M_{\rm{BH}}$) predicted by their simulation-based inference model (cranmer2020) can be biased by up to 0.5 dex when the test light curves are five times noisier than those used for training.
In this paper, we present EzTaoX—a scalable and flexible AGN light curve modeling tool that leverages LSST’s multiband observations to simultaneously infer the properties of the underlying stochastic variability and the lags between photometric bands (eztaox). The name of EzTaoX combines three elements: “Ez", a homophone of easy; “Tao" referencing the Chinese philosophical and religious tradition of Taoism, commonly translated as “the Way"; and “X" denoting that the code is implemented in JAX—a high-performance numerical computing library developed by Google. EzTaoX achieves a speed increase of $\sim 10^{2}-10^{4}\times$ on CPUs compared to existing tools with similar capabilities (e.g., JAVELIN), making it feasible to perform both stochastic variability characterization and interband continuum lag measurements for all AGNs discovered by LSST. Furthermore, EzTaoX’s multiband capability enables robust constraints on stochastic variability timescales (e.g., $\tau_{\rm DRW}$) down to one hour with the LSST WFD cadence, facilitating the identification of low-mass AGNs—such as those residing in dwarf galaxies—which are expected to exhibit characteristic $\tau_{\rm DRW}$ ranging from hours to days (burke2021; Wang2023a).
@@ 2 Light Curve Modeling Method & Implementation
EzTaoX models AGN UV/optical stochastic variability as a Gaussian process (GP), a method commonly adopted in AGN variability analysis (e.g., Press1992; kelly2009; kozlowski2010; Wilkins2019; Griffiths2021; yu2022b; stone2022; Zhang2023; McLaughlin2024). Since the flux light curves of accreting compact objects typically follow a lognormal distribution (e.g., uttley2001; gaskell2004; Giebels2009; macleod2012; decicco2022), we emphasize that GP modeling is best applied to magnitude light curves, which follow a Gaussian distribution. Moreover, Gurpide2025 demonstrated that even when the light curves follow a lognormal distribution, the simulated GP parameters can be recovered without significant bias.
Given a single band AGN light curve, a GP treats the intrinsic brightness of the AGN at each point in time as a Gaussian random variable and captures the temporal correlations between observations through a kernel function $k(t_{i},t_{j})$, which defines the covariance between the Gaussian variables at any two timestamps $t_{i}$ and $t_{j}$ (Rasmussen2006; aigrain2023).
@@
where $k_{\rm latent}$ denotes the kernel of the underlying latent GP. $S_{1}$ and $S_{2}$ are the variability-amplitude scaling factors for the corresponding photometric bands, and $\tau_{\rm lag}$ represents the relative time lag between them. The variability amplitude in each band is obtained by multiplying the latent GP amplitude by the corresponding scaling factor. This class of kernel functions, as expressed in Equation 5, has seen broad application in astronomical time series analysis (e.g., gordon2020; Villar2021). To evaluate the likelihood of the multiband model for a given light curve, we shift the light curve in different bands by their trial $\tau_{\rm lag}$ relative to the reference band, and compute Equation 2 using a kernel that assumes no intrinsic lag between bands (i.e., setting $\tau_{\rm lag}=0$ in Equation 5). In this way, the GP parameters and the interband lags are fitted jointly.
@@
EzTaoX is built on top of tinygp [^2] https://github.com/dfm/tinygp—a GP modeling framework developed using JAX. [^3] https://github.com/jax-ml/jax JAX is a high-performance numerical computing library that combines automatic differentiation, just-in-time (JIT) compilation, and hardware acceleration to empower scalable machine learning and scientific computing. tinygp leverages these features of JAX and re-implements the novel celerite algorithm introduced by foreman-mackey2017 in JAX. When a kernel function can be expressed as a mixture of exponentials, the celerite algorithm can evaluate the GP likelihood function with $\mathcal{O}(N)$ computational scaling, a significant efficiency enhancement over the standard $\mathcal{O}(N^{3})$ scaling associated with direct likelihood evaluation (foreman-mackey2017). EzTaoX takes advantage of the JAX-based implementation of the celerite algorithm to provide a scalable solution for multiband light curve modeling of AGNs.
@@
PSD shapes exhibited by the light curves of accreting black holes can be effectively modeled with celerite GP kernels. For example, the continuous-time autoregressive moving-average (CARMA) process—commonly used as a more flexible alternative to the DRW process—has an equivalent celerite implementation (rouxalet2002; kelly2014; kasliwal2017; moreno2019; yu2022b). A CARMA($p$, $q$) process combines an autoregressive process of order $p$ with a moving-average process of order $q$, and its PSD can be expressed as a weighted sum of modified Lorentzians. The DRW process is a special case of a CARMA process with $p=1$ and $q=0$. For a detailed description of CARMA modeling, we refer the reader to kelly2014. More recently, Lefkir2025 introduced an algorithm for approximating arbitrary broken PL PSDs using a sum of multiple celerite kernels, thereby enhancing the interpretability and flexibility of GP-based modeling for AGN light curves. Moreover, celerite-based kernels can be freely combined through addition and multiplication to model more complex covariance structures, while retaining the $\mathcal{O}(N)$ computational scaling. All such kernels can be used for simultaneous multiband variability characterization and interband lag measurements.